dice with four faces was rolled
x 1 2 3 4
f(x) 0.16 0.19 0.21 0.44
What is the probability that at least a 3 is rolled?

Answer by stanbon(75887)

dice with four faces was rolled
x 1 2 3 4
f(x) 0.16 0.19 0.21 0.44
What is the probability that at least a 3 is rolled?
Ans: (x = 3 or 4) = P(x=3)+P(x=4) = 0.21+0.44 = 0.65
-------------------
Cheers,
Stan H.
===================